# prompt[Prompt][List of Prompts]

### v3

@@ 38       $\mathcal{A}_{\mathcal{V}}\leftarrow(Q,q_{0},\delta,\lambda)$
Object Existence: G (“record_player” & “lamp”)
Object Action Alignment: G((“record_player_spins”&“lamp_glows”))
Spatial Relationship: G(“vinyl_on_record_player” & “light_glowing_from_nearby_lamp”)
Overall Consistency: G (“record_player_spinning_a_vinyl” & “light_glowing_softly_from_a_nearby_lamp”)
@@
Object Existence: (“drone”) $\mathsf{U}$ (“waypoint”)
Object Action Alignment: (“drone_hovers_in_air” $\mathsf{U}$ “drone_reaches_next_waypoint”) & F (“drone_continues_to_fly”)
Spatial Relationship: (“drone_in_air”) $\mathsf{U}$ (“drone_reaches_waypoint”)
Overall Consistency: (“drone_hovering_in_air” $\mathsf{U}$ “drone_reaches_next_waypoint”) & F (“drone_continues_to_fly”)
@@
Object Existence: (“lightbulb”) $\mathsf{U}$ “switch”
Object Action Alignment: ((“lightbulb_flickers” $\mathsf{U}$ “switch_turned_off”)) & F(“room_darkens”)
Spatial Relationship: (“lightbulb_flickering” $\mathsf{U}$ “switch_is_turned_off”)
Overall Consistency: (“lightbulb_flickering_intermittently” $\mathsf{U}$ “switch_is_turned_off”) & F “room_is_cast_into_darkness”
@@
Object Existence: “vehicle” $\mathsf{U}$ “stop_sign”
Object Action Alignment: (“vehicle_moves” $\mathsf{U}$ “vehicle_reaches_stop_sign”)
Spatial Relationship: (“vehicle_moving_forward” $\mathsf{U}$ “vehicle_at_stop_sign”)
Overall Consistency: (“vehicle_moving_forward” $\mathsf{U}$ “vehicle_reaches_a_stop_sign”)
@@
Object Existence: (“motorcycle” & “bus”)
Object Action Alignment: (“motorcycle_revs”&“bus_pulls_up”)
Spatial Relationship: (“motorcycle_revving” & “bus_pulling_up_beside”)
Overall Consistency: (“motorcycle_revving” & “bus_pulling_up_beside”)
@@
Object Existence: (“traffic_light”) $\mathsf{U}$ “pedestrians”
Object Action Alignment: ((“traffic_light_is_red” $\mathsf{U}$ “pedestrians_finish_crossing”)) & F(“traffic_light_turns_green”)
Spatial Relationship: (“traffic_light_turning_red” $\mathsf{U}$ “pedestrians_finish_crossing”)
Overall Consistency: (“traffic_light_turning_red” $\mathsf{U}$ “pedestrians_finish_crossing”) & F “traffic_light_shifts_to_green”
@@
Object Existence: G (“vehicle” & “driver”)
Object Action Alignment: G((“electric_vehicle_charges”&“driver_reads”))
Spatial Relationship: G(“electric_vehicle_charging_at_station” & “driver_reading_book_nearby”)
Overall Consistency: G (“electric_vehicle_charging_at_the_station” & “driver_reading_a_book_nearby”)
@@
Object Existence: (“car” & “streets”) $\mathsf{U}$ (“construction_zone” & “path”)
Object Action Alignment: (“car_drives_through_city_streets” $\mathsf{U}$ “car_encounters_construction_zone”) & F (“car_reroutes_to_alternate_path”)
Spatial Relationship: (“car_on_city_streets”) $\mathsf{U}$ (“car_at_construction_zone”)
Overall Consistency: (“car_driving_through_city_streets” $\mathsf{U}$ “it_encounters_a_construction_zone”) & F (“it_reroutes_to_an_alternate_path”)
Table 7: NeuS-V Prompt Suite (continued): Illustrative prompts and their detailed specifications (across all four modes) for varying complexities within the “Object Interactions” and “Driving Data” themes.

### v5

@@ 38       $\mathcal{A}_{\mathcal{V}}\leftarrow(Q,q_{0},\delta,\lambda)$
Object Existence: G (“record_player" & “lamp")
Object Action Alignment: G((“record_player_spins"&“lamp_glows"))
Spatial Relationship: G(“vinyl_on_record_player" & “light_glowing_from_nearby_lamp")
Overall Consistency: G (“record_player_spinning_a_vinyl" & “light_glowing_softly_from_a_nearby_lamp")
@@
Object Existence: (“drone") $\mathsf{U}$ (“waypoint")
Object Action Alignment: (“drone_hovers_in_air" $\mathsf{U}$ “drone_reaches_next_waypoint") & F (“drone_continues_to_fly")
Spatial Relationship: (“drone_in_air") $\mathsf{U}$ (“drone_reaches_waypoint")
Overall Consistency: (“drone_hovering_in_air" $\mathsf{U}$ “drone_reaches_next_waypoint") & F (“drone_continues_to_fly")
@@
Object Existence: (“lightbulb") $\mathsf{U}$ “switch"
Object Action Alignment: ((“lightbulb_flickers" $\mathsf{U}$ “switch_turned_off")) & F(“room_darkens")
Spatial Relationship: (“lightbulb_flickering" $\mathsf{U}$ “switch_is_turned_off")
Overall Consistency: (“lightbulb_flickering_intermittently" $\mathsf{U}$ “switch_is_turned_off") & F “room_is_cast_into_darkness"
@@
Object Existence: “vehicle" $\mathsf{U}$ “stop_sign"
Object Action Alignment: (“vehicle_moves" $\mathsf{U}$ “vehicle_reaches_stop_sign")
Spatial Relationship: (“vehicle_moving_forward" $\mathsf{U}$ “vehicle_at_stop_sign")
Overall Consistency: (“vehicle_moving_forward" $\mathsf{U}$ “vehicle_reaches_a_stop_sign")
@@
Object Existence: (“motorcycle" & “bus")
Object Action Alignment: (“motorcycle_revs"&“bus_pulls_up")
Spatial Relationship: (“motorcycle_revving" & “bus_pulling_up_beside")
Overall Consistency: (“motorcycle_revving" & “bus_pulling_up_beside")
@@
Object Existence: (“traffic_light") $\mathsf{U}$ “pedestrians"
Object Action Alignment: ((“traffic_light_is_red" $\mathsf{U}$ “pedestrians_finish_crossing")) & F(“traffic_light_turns_green")
Spatial Relationship: (“traffic_light_turning_red" $\mathsf{U}$ “pedestrians_finish_crossing")
Overall Consistency: (“traffic_light_turning_red" $\mathsf{U}$ “pedestrians_finish_crossing") & F “traffic_light_shifts_to_green"
@@
Object Existence: G (“vehicle" & “driver")
Object Action Alignment: G((“electric_vehicle_charges"&“driver_reads"))
Spatial Relationship: G(“electric_vehicle_charging_at_station" & “driver_reading_book_nearby")
Overall Consistency: G (“electric_vehicle_charging_at_the_station" & “driver_reading_a_book_nearby")
@@
Object Existence: (“car" & “streets") $\mathsf{U}$ (“construction_zone" & “path")
Object Action Alignment: (“car_drives_through_city_streets" $\mathsf{U}$ “car_encounters_construction_zone") & F (“car_reroutes_to_alternate_path")
Spatial Relationship: (“car_on_city_streets") $\mathsf{U}$ (“car_at_construction_zone")
Overall Consistency: (“car_driving_through_city_streets" $\mathsf{U}$ “it_encounters_a_construction_zone") & F (“it_reroutes_to_an_alternate_path")
Table 7: NeuS-V Prompt Suite (continued): Illustrative prompts and their detailed specifications (across all four modes) for varying complexities within the “Object Interactions" and “Driving Data" themes.